# Dynamics of Spontaneous Topic Changes in Next Token Prediction with Self-Attention

### v3

@@ 4 Attention within Mixed Topics
Let’s first understand the way in which attention models assign priority to tokens within mixed-topic setting. For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend the results on multiple topics. Notice the self-attention embedding output is a linear combination of $\bf X$ given by $\mathbb{S}({\bf XW\bar{x}})$. The embeddings in $\bf X$ corresponding to the highest entries in $\mathbb{S}({\bf XW\bar{x}})$ will receive higher priority to predict the next token, therefore we can hypothesize that models in which $\mathbb{S}({\bf XW\bar{x}})$ are ordered in a similar way will predict similar next tokens. This idea leads to our first main result which considers this situation within a mixed topics setting:
@@
This implies that for an input sequence $\mathbf{X}$, a model trained in a mixed-topic setting will maintain the priority of the topic to which $\mathbf{X}$ belongs. Consequently, the attention will be allocated in the same order as if the model had been trained exclusively on the original topic of $\mathbf{X}$. For the first input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$ from $\mathbb{T}_{a}$, as shown in Figure 3 (right), the predicted next token $\hat{y}_{{\bf w}_{ab}}$ is ${\bf e}_{5}$ and the highest probability SCC in mixed-topics is $\widehat{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$. Since $\hat{y}_{{\bf w}_{ab}}$ belongs to ${\widehat{\mathcal{G}}}^{(4)}_{ab}({\bf X})$, ${\bf W}_{ab}$ for input sequence $\bf X$ is considered as topic continuity, based on the Definition 3.
The only assumption about ${\bf X}$ on Theorem 3 is that it belongs to $\mathbb{T}_{a}$. However, if $\mathbf{X}$ belongs to $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, the priority will be preserved within both topics. Additionally, strict equality in the attention priority holds, but strict inequalities may not, as the union of their TPGs can form new SCCs, like $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$ in $\mathcal{G}^{(4)}_{ab}$ (left side of Figure 3), disrupting strict priority (see details in Appendix C).
Figure 3: Depiction of each scenario in next token prediction. Left: Taking the last token ${\bf e}_{4}$ as an example, $\mathcal{G}^{(4)}_{ab}$ for $\mathbb{T}_{ab}$ is formed by the union of $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$. Right: For each input sequence belonging to $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens, denoted as $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$ (orange). $\widehat{\mathcal{G}}^{(4)}_{ab}$ and $\widehat{\mathcal{G}}^{(4)}_{a}$ represent the highest probability SCCs (Definition 1) in mixed-topics setting and in $\mathbb{T}_{a}$, respectively. There are three scenarios, topic continuity (Definition 3), ambiguous sequence (Definition 4), and change of topic (Definition 5). The numeric details for each scenario are provided in Appendix D.5.
@@ 5 Explaining Topic Shifts
The formation of new SCCs when combining datasets suggests that the highest priority SCC for some input sequences may increase in size in this new setting. This also suggests that topic shifts may arise from ambiguity within an input sequence rather than a straightforward change in topic. In our analogy on the oracle, gaining knowledge of both Topic A and Topic B might cause a conversation to be naturally followed within Topic A or also outside Topic A. We introduce the following definition to characterize this phenomenon:
@@
    If $p>\max_{x\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$,
    then $\lim_{T\to\infty}{\mathbb{P}(\tilde{\bf W}_{ab}\text{ changes topic }\mathbb{T}_{a}\text{ for }{\bf X}^{\prime})}=0$.
@@
There are two implications of this theorem. First, as the input sequence length increases sufficiently, the likelihood of topic changes vanishes. Second, notice that increasing $l$ increases the probability of overlap between topics, and it also decreases the probability of satisfying the necessary conditions for a topic change, effectively creating a bound on the frequency of topic changes. In practice, consider the oracle analogy: if the oracle is following a sufficiently long conversation on a specific topic, it becomes exceedingly unlikely to shift topics. Similarly, as topics A and B become more interconnected, this increased ambiguity does not lead to more topic changes; rather, it may reduce their occurrence. This contrasts with human cognition, where longer conversations and greater inter-connectivity of knowledge increase the likelihood of spontaneous shifts.
To illustrate Theorem 5 with simulations, we generate embeddings with $K=10$ and $d=16$. We approximate $\tilde{\mathbf{W}}_{a}$ and $\tilde{\mathbf{W}}_{ab}$ as the results obtained after $\tau=8000$ iterations of Algo-GD. We quantify the proportion of test sequences in which $\tilde{{\bf W}}_{ab}$ keeps $\mathbb{T}_{a}$ (keep topic), proportion of ambiguous sequences in $\mathbb{T}_{ab}$ (ambiguity) and proportion of times in which $\tilde{{\bf W}}_{ab}$ changes topic (change topic). First, we explore the effect of longer sequences by varying the length $T$ of the test sequences $\bf Z$. We increase $T$ from 4 to 512. Figure 5 illustrates how the proportion of change topic decreases as $T$ increases. Second, we investigate the effect of overlapping topics by generating TPGs with an increased number of edges $L$. Intuitively, a higher $L$ results in an increase $l$ and a greater overlap between TPGs of different topics. We vary $L$ from 4 to 18. Figure 5 demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic doesn’t increase. These two findings contrast with expectations derived from human cognition but align with the result of Theorem 5. Lastly, among the 85000 test sequences generated for these experiments, 99.98% satisfy Theorem 4 (i.e., topic changes occur when a low-priority token appears more frequently than high-priority tokens). The remaining 0.02% mismatched cases were solely due to minor approximation discrepancies in the attention softmax. These results validate Theorem 4 (See simulation details in Appendix D).
(a) Input length.
@@
Figure 4: The proportion of topic continuity, ambiguity, and topic change as (a) input length and (b) topic ambiguity increase.
(a) Input length.
@@
Figure 5: Cosine similarity between continuations generated with single-topic and mixed-topic knowledge as (a) input length and (b) topic ambiguity vary.
6 Experiments in Frontier LLMs
To prove Theorem 5 we work within the simplified, single-layer self-attention model of Li et al., 2024b . Although this abstraction omits many hallmarks of contemporary LLMs (deep stacks of attention blocks, alternative cost functions, and other training heuristics), it offers a mathematically tractable setting that lets us derive interesting mathematical results. These results, in turn, can be used to understand how cutting-edge LLMs behave in terms of spontaneous topic changes. We empirically investigate such behavior on four frontier models: GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3.
Experimental setup. For two distinct papers A and B, and an input prompt ($\bf X)$ from paper A, we consider a measure of topic continuity as the cosine similarity between the embeddings of the texts generated when the LLM has contextual knowledge solely from paper A ($\hat{y}_{\bf{W}_{a}}$) and when the LLM has contextual knowledge from both paper A and B ($\hat{y}_{\bf{W}_{ab}}$). We consider this cosine similarity as an empirical proxy for our definition of topic continuity (Definition 3): therefore the larger the similarity, the smaller the chance that the model has lead to a change of topic. This proxy suggests two testable consequences which become the empirical counterpart of our Theorem 5: Cosine similarity is expected to increase with the length of the input prompt, and is not expected to decrease when increasing ambiguity between paper A and paper B.
To more closely align with our theoretical framework, where a model gains knowledge of topic A and incrementally gains knowledge of topic B, we implement a Retrieval-Augmented Generation (RAG) approach, retrieving information exclusively from paper A or jointly from papers A and B (Von Oswald et al.,, 2023). Based on input prompt, we retrieve the top 3 most relevant excerpts from paper A or B to form the contextual knowledge set A or set B. The combined contextual knowledge set is simply the union of sets A and B. We add set A to the input prompt to obtain the generated text with sole knowledge of paper A ($\hat{y}_{\bf{W}_{a}}$), and we add the combined set to the input prompt to obtain the generated text with combined knowledge of paper A and B ($\hat{y}_{\bf{W}_{ab}}$). To closely follow our greedy decoding approach in our theoretical framework, we set the temperature parameter to 0 for all LLMs.
We designate each paper as paper A, and we randomly select 5 different papers from the remaining 99 papers as distinct paper B. For each input segment, we calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across these five pairs of paper A and paper B, using each LLM. The results for each LLM are averaged over the 100 papers (See more experimental details in Appendix E).
Experiment 1: Impact of the input length. We use the first $10,30,\ldots,150$ words from each paper A’s abstract as the input prompt. Figure 5 plots, for each LLM, the average cosine similarity as a function of the input length; the shaded bands indicate 95% confidence intervals. Across all models, similarity tends to increase with prompt length, aligning with the behavior predicted by Theorem 5.
Experiment 2: Impact of the topic ambiguity. We fix the input prompt length to the first $80$ words for each paper A’s abstract. We quantify paper ambiguity by the average similarity among each paper’s keywords: higher keyword similarity signifies lower paper ambiguity because topic is more specific. We partitioned the papers into six equal-width bins along this ambiguity spectrum. Figure 5 summarises the results: each boxplot shows the distribution of cosine similarities within an ambiguity bin (x-axis ordered from least to most ambiguous). Across all LLMs the median similarity does not seem to decrease, in agreement with the prediction of Theorem 5.
Taken together, the two experiments provide preliminary empirical support for Theorem 5, showing that its prediction, derived for a single-layer self-attention toy model, can be extended to today’s deep, multi-layer LLMs. Crucially, an important divergence between machine and human cognition persists in these frontier models: neither longer prompts nor greater topic ambiguity appreciably raise the likelihood of an spontaneous topic change.
@@ 7 Related Work
Training and generalization of Transformer. (1) Properties of Softmax. The self-attention mechanism employs the softmax function to selectively emphasize different parts of the input. Gu et al., (2024), Goodfellow et al., (2016), and Deng et al., (2023) underscore the pivotal role of the softmax function in shaping attention distributions, influencing how models process and prioritize information within input sequences. Bombari and Mondelli, (2024) examined the word sensitivity of attention layers, revealing that softmax-based attention-layers are adept at capturing the significance of individual words. However, recent work has also pointed out limitations of the softmax function (Saratchandran et al.,, 2024; Deng et al.,, 2023). (2) Optimization in attention-based models. Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ; Li et al., 2024b, ; Julistiono et al.,, 2024). (3) Chain-of-Thought (CoT) and In-Context Learning (ICL). Moreover, transformers exhibit remarkable abilities in generalization through ICL, where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022). CoT prompting (Wei et al.,, 2022; Zhou et al.,, 2023; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers. (4) Improvement efficiency of transformers. Recent advancements aim to improve the computational efficiency of transformers (Kitaev et al.,, 2020; Choromanski et al.,, 2021; Sukhbaatar et al.,, 2019; Wang et al.,, 2020), ensuring their viability for large-scale deployment while maintaining or enhancing their representational capabilities.
@@
Next token prediction in LLMs. (1) Theoretical and architectural innovations. Shannon, (1951)’s foundational work laid the groundwork for estimating the predictability of natural language sequences, providing a basis for subsequent advances in language modeling. Recent studies have expanded our understanding of how LLMs anticipate future tokens from internal hidden states, offering valuable insights into the efficiency and effectiveness of transformer-based architectures (He and Su,, 2024; Pal et al.,, 2023; Shlegeris et al.,, 2024). Despite their impressive predictive capabilities, these models face fundamental limitations. For instance, Bachmann and Nagarajan, (2024) highlights the shortcomings of teacher-forced training, emphasizing how this approach can fail and suggesting strategies to improve model robustness. (2) Efficiency and Optimization. Goyal et al., (2024) introduces a novel method that incorporates a deliberate computation step before output generation, enhancing reasoning capabilities. Additionally, Gloeckle et al., (2024) advocates for multi-token prediction, which significantly improves both efficiency and speed.
@@
Self-Attention and topic dynamics. Advancements in self-attention research have deepened our understanding of how transformers handle evolving semantic contexts. Prior work has explored diverse aspects of topic modeling, such as dynamic topic structures (Miyamoto et al.,, 2023), hierarchical relationships (Lin et al.,, 2024), topic-aware attention mechanisms (Panwar et al.,, 2021), and the mechanistic underpinnings of topic representation (Li et al.,, 2023). While these studies provide insights into managing static and hierarchical topic structures, our work focuses on the topic changes with the given input sequences from a specific topic.

### v4

@@ 4 Attention within mixed topics
Let’s first understand how attention models assign priority to tokens within mixed-topic setting. For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend the results on multiple topics. Notice the self-attention embedding output is a linear combination of $\bf X$ given by $\mathbb{S}({\bf XW\bar{x}})$. The embeddings in $\bf X$ corresponding to the highest entries in $\mathbb{S}({\bf XW\bar{x}})$ will receive higher priority to predict the next token, therefore we can hypothesize that models in which $\mathbb{S}({\bf XW\bar{x}})$ are ordered in a similar way will predict similar next tokens. This idea leads to our first main result which considers this situation within a mixed-topic setting:
@@
This implies that for an input sequence $\mathbf{X}$, a model trained in a mixed-topic setting will maintain the priority of the topic to which $\mathbf{X}$ belongs. Consequently, the attention will be allocated in the same order as if the model had been trained exclusively on the original topic of $\mathbf{X}$. For the first input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$ from $\mathbb{T}_{a}$, as shown in Figure 3 (right), the predicted next token $\hat{y}_{{\bf w}_{ab}}$ is ${\bf e}_{5}$ and the highest probability SCC in mixed topics is $\widehat{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$. Since $\hat{y}_{{\bf w}_{ab}}$ belongs to ${\widehat{\mathcal{G}}}^{(4)}_{ab}({\bf X})$, ${\bf W}_{ab}$ for input sequence $\bf X$ is considered as topic continuity, based on the Definition 3.
The only assumption about $\bf X$} on Theorem 2 is that it belongs to $\mathbb{T}_{a}$. However, if $\mathbf{X}$ belongs to $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, the priority will be preserved within both topics. Additionally, strict equality in the attention priority holds, but strict inequalities may not, as the union of their TPGs can form new SCCs. As illustrated on the left of Figure 3, $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ denote the TPGs corresponding to the last input token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. In $\mathcal{G}^{(4)}_{a}$, the token priority is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. In contrast, in $\mathcal{G}^{(4)}_{ab}$ for the mixed topics, the priority order is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}={\bf e}_{4}$. The equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, whereas the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is relaxed to ${\bf e}_{2}={\bf e}_{4}$ in mixed topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
Figure 3: Depiction of each scenario in next token prediction. Left: Taking the last token ${\bf e}_{4}$ as an example, $\mathcal{G}^{(4)}_{ab}$ for $\mathbb{T}_{ab}$ is formed by the union of $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$. The direction of edge is from output to input and the dotted square denotes the strongly-connected components (SCC) in which tokens have equal priority. Right: For each input sequence belonging to $\mathbb{T}_{a}$, we use a self-attention model trained on $\text{DSET}_{a}$ and another model trained on the mixed-topic dataset $\text{DSET}_{ab}$ to predict the next tokens, denoted as $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$, respectively. $\widehat{\mathcal{G}}^{(4)}_{ab}$ and $\widehat{\mathcal{G}}^{(4)}_{a}$ represent the highest probability SCCs (Definition 1) in mixed-topic setting and in $\mathbb{T}_{a}$, respectively. There are three scenarios, topic continuity (Definition 3), ambiguous sequence (Definition 4), and change of topic (Definition 5). The numeric details for each scenario are provided in Appendix C.5.
@@ 5 Explaining topic shifts
The formation of new SCCs when combining datasets suggests that the highest priority SCC for some input sequences may increase in size in this new setting. This also suggests that topic shifts may arise from ambiguity within an input sequence rather than a straightforward change in topic. In our oracle analogy, gaining knowledge of both Topic A and Topic B might cause a conversation to be naturally followed within Topic A or also outside Topic A. We introduce the following definition to characterize this phenomenon:
@@
    If $p>\max_{x\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$, then $\lim_{T\to\infty}{\mathbb{P}(\tilde{\bf W}_{ab}\text{ changes topic }\mathbb{T}_{a}\text{ for }{\bf X}^{\prime})}=0$.
@@
There are two implications of this theorem. First, as the input sequence length increases sufficiently, the likelihood of topic changes vanishes. Second, increasing $l$ raises the probability of overlap between topics and reduces the probability of satisfying the necessary conditions for a topic change, effectively creating a bound on the frequency of topic changes. In practice, consider the oracle analogy: if the oracle is following a sufficiently long conversation on a specific topic, it becomes exceedingly unlikely to shift topics. Similarly, as topics A and B become more interconnected, this increased ambiguity does not lead to more topic changes; rather, it may reduce their occurrence. This contrasts with human cognition, where longer conversations and greater inter-connectivity of knowledge increase the likelihood of spontaneous topic changes.
To illustrate Theorem 4 through simulations, we generate embeddings with $K=10$ and $d=16$. We approximate $\tilde{\mathbf{W}}_{a}$ and $\tilde{\mathbf{W}}_{ab}$ as the results obtained after $\tau=8000$ iterations of Algo-GD. We quantify the proportion of test sequences in which $\tilde{{\bf W}}_{ab}$ keeps $\mathbb{T}_{a}$ (keep topic), proportion of ambiguous sequences in $\mathbb{T}_{ab}$ (ambiguity) and proportion in which $\tilde{{\bf W}}_{ab}$ changes topic (change topic). First, we explore the effect of longer sequences by varying the length $T$ of the test sequences $\bf Z$. We increase $T$ from 4 to 512. Figure 4 illustrates how the proportion of change topic decreases as $T$ increases. Second, we investigate the effect of topic overlap with an increasing number of edges $L$. Intuitively, a higher $L$ results in an increase $l$ and a greater overlap between TPGs of different topics. We vary $L$ from 4 to 18. Figure 4 demonstrates that as $L$ increases, ambiguity increases, while the proportion of change topic doesn’t increase. These two findings contrast with expectations derived from human cognition but align with the result of Theorem 4. Lastly, among the 85,000 test sequences generated for these experiments, 99.98% satisfy Theorem 3 (i.e., topic changes occur when a low-priority token appears more frequently than high-priority tokens). The remaining 0.02% mismatched cases are solely due to minor approximation discrepancies in the attention softmax. These results validate Theorem 3 (see simulation details in Appendix C).
(a) Length of sequences.
@@
Figure 4: The proportion of topic continuity, ambiguous sequence, and change of topic as (a) input length and (b) topic ambiguity increase.
6 Experiments in frontier LLMs
To prove Theorem 4 we work within the simplified, single-layer self-attention model of li2024mechanics. Although this abstraction omits many hallmarks of contemporary LLMs (deep stacks of attention blocks, alternative cost functions, and other training heuristics), it offers a mathematically tractable setting that lets us derive interesting mathematical results. These results, in turn, can be used to understand how cutting-edge LLMs behave in terms of spontaneous topic changes. We empirically investigate such behavior on four frontier models: GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3.
Experimental setup. For two distinct papers A and B, and an input prompt ($\bf X)$ from paper A, we consider a measure of topic continuity as the cosine similarity between the embeddings of the texts generated when the LLM has contextual knowledge solely from paper A ($\hat{y}_{\bf{W}_{a}}$) and when the LLM has contextual knowledge from both paper A and B ($\hat{y}_{\bf{W}_{ab}}$). We treat this cosine similarity as an empirical proxy for our formal definition of topic continuity (Definition 3): therefore the larger the similarity, the smaller the chance that the model has led to a change of topic. This proxy suggests two testable consequences which become the empirical counterpart of our Theorem 4: (1) cosine similarity is expected to increase with the length of the input prompt, and (2) it is not expected to decrease with increasing ambiguity in paper A and paper B.
To more closely align with our theoretical framework, where a model gains knowledge of topic A and incrementally gains knowledge of topic B, we implement a Retrieval-Augmented Generation (RAG) approach, retrieving information exclusively from paper A or jointly from papers A and B [von2023transformers]. Based on the input prompt, we retrieve the top 3 most relevant excerpts from paper A or paper B to form the contextual knowledge set A or set B. The combined contextual knowledge set is simply the union of sets A and B. We add set A to the input prompt to obtain the generated text with sole knowledge of paper A ($\hat{y}_{\bf{W}_{a}}$), and we add the combined set to the input prompt to obtain the generated text with combined knowledge of paper A and B ($\hat{y}_{\bf{W}_{ab}}$). To closely follow our greedy decoding approach in our theoretical framework, we set the temperature parameter to 0 for all LLMs.
We designate each paper as paper A and randomly select 5 different papers from the remaining 99 papers as distinct paper B. For each input segment, we calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across these five pairs of paper A and paper B, using each LLM. The results for each LLM are averaged over all 100 papers. See additional experimental details in Appendix D.
Experiment 1: Impact of input length. We use the first $10,30,\ldots,150$ words from each paper A’s abstract as the input prompt. Figure 5 shows, for each LLM, the average cosine similarity as a function of input length; shaded bands indicate 95% confidence intervals. Across all models, similarity tends to increase with input length, aligning with the behavior predicted by Theorem 4. Appendix D.3.1 presents an additional experiment in which we extend the input length to 1210 words extracted from each paper’s introduction; the results further support our conclusions.
(a) Length of sequences.
@@ 6 Experiments in frontier LLMs
Figure 5: Similarity between continuations generated with single-topic and mixed-topic knowledge as (a) input length and (b) topic ambiguity increase.
Experiment 2: Impact of topic ambiguity. We fix the input prompt length to the first $80$ words of each paper A’s abstract. We quantify topic ambiguity by the average similarity among each paper A’s keywords: lower keyword similarity signifies higher probability of overlap between paper A and other papers, consistent with our setup in Theorem 4. We partition the papers into five equal-width bins along this ambiguity spectrum. Figure 5 summarizes the results: each boxplot shows the distribution of cosine similarities within an ambiguity bin, with the x-axis ordered from least to most ambiguous. Across all LLMs the median similarity does not seem to decrease, in agreement with the prediction of Theorem 4. In Appendix D.3.2, we present an additional experiment using an alternative ambiguity measure based on cross-paper keyword similarity, yielding results consistent with our conclusions.
Taken together, the two experiments provide preliminary empirical support for Theorem 4, showing that its prediction, derived from a single-layer self-attention toy model, can be extended to today’s deep, multi-layer LLMs. Crucially, an important divergence between machine and human cognition persists in these frontier models: neither longer prompts nor greater topic ambiguity appreciably increases the likelihood of a spontaneous topic change.
@@ 7 Related work
Training and generalization of Transformer. (1) Properties of Softmax. The self-attention mechanism employs the softmax function to selectively emphasize different parts of the input. gu2024exploring, Goodfellow-et-al-2016, and deng2023superiority underscore the pivotal role of the softmax function in shaping attention distributions, influencing how models process and prioritize information within input sequences. bombari2024towards examined the word sensitivity of attention layers, revealing that softmax-based attention layers are adept at capturing the significance of individual words. However, recent work has also pointed out limitations of the softmax function [saratchandran2024rethinking, deng2023superiority]. (2) Optimization in attention-based models. Additionally, recent research interprets Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space [tarzanagh2023transformers, ataee2023max, li2024mechanics, julistiono2024optimizing]. (3) Chain-of-Thought (CoT) and In-Context Learning (ICL). Moreover, transformers exhibit remarkable abilities in generalization through ICL, where models effectively learn from contextual cues during inference [brown2020language, xie2021explanation, olsson2022context]. CoT prompting [wei2022chain, zhou2022least, pmlr-v202-shao23a, li2024training] enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers. (4) Improvement efficiency of transformers. Recent advancements aim to improve the computational efficiency of transformers [Kitaev2020Reformer, choromanski2021rethinking, sukhbaatar-etal-2019-adaptive, wang2020linformer], ensuring their viability for large-scale deployment while maintaining or enhancing their representational capabilities.
@@
Next token prediction in LLMs. (1) Theoretical and architectural innovations. shannon1950prediction’s foundational work laid the groundwork for estimating the predictability of natural language sequences, providing a basis for subsequent advances in language modeling. Recent studies have expanded our understanding of how LLMs anticipate future tokens from internal hidden states, offering valuable insights into the efficiency and effectiveness of Transformer-based architectures [he2024law, pal2023future, shlegeris2024languagemodelsbetterhumans]. Despite their impressive predictive capabilities, these models face fundamental limitations. For instance, pmlr-v235-bachmann24a highlights the shortcomings of teacher-forced training, emphasizing how this approach can fail and suggesting strategies to improve model robustness. (2) Efficiency and Optimization. goyal2024think introduces a novel method that incorporates a deliberate computation step before output generation, enhancing reasoning capabilities. Additionally, gloeckle2024better advocates for multi-token prediction, which significantly improves both efficiency and speed.
@@
Self-Attention and topic dynamics. Advancements in self-attention research have deepened our understanding of how transformers handle evolving semantic contexts. Prior work has explored diverse aspects of topic modeling, such as dynamic topic structures [miyamoto-etal-2023-dynamic], hierarchical relationships [lin-etal-2024-hierarchical], topic-aware attention mechanisms [panwar-etal-2021-tan], and the mechanistic underpinnings of topic representation [li2023transformers]. While these studies provide insights into managing static and hierarchical topic structures, our work focuses on the topic changes with the given input sequences from a specific topic.